# Optimal Multi-Distribution Learning

### v2

@@ C.3 Statement and proof of Lemma 29
For $1\leq j\leq j_{\mathrm{max}}:=\left\lfloor 2\log_{2}(k)+1\right\rfloor$, we let
$\tau_{j}:=\min\{\tau|t_{1}\leq\tau\leq t_{2}-1,\,-v^{\tau}\leq 2^{-(j-1)}\}.$
We also define $\tau_{j_{\mathrm{max}+1}}=t_{2}$. By definition, we know that $\tau_{1}=t_{1}$ and $\tau_{j_{2}}\geq\tau_{j_{1}}$ for $j_{2}\geq j_{1}$. Let $\mathcal{I}$ be the index set of this segment. Let $y_{j}:=\sum_{i\in\mathcal{I}}w_{i}^{\tau_{j}}$.
We then claim that there exists $1\leq\tilde{j}\leq j_{\mathrm{max}}$ such that
$\log\left(\frac{y_{\tilde{j}+1}}{y_{\tilde{j}}}\right)$ $\geq\frac{x}{2\log_{2}(k)+1}.$ (94)
Suppose that none of $1\leq\tilde{j}\leq j_{\mathrm{max}}$ satisfies (94). Then for any $j$, it holds that $\log\left(\frac{y_{j+1}}{y_{j}}\right)\leq\frac{x}{2\log_{2}(k)+1}$, which implies that $y_{j}\geq y_{j+1}\exp\left(\frac{x}{2\log_{2}(k)+1}\right)$. As a result, we have that $y_{1}\geq y_{j_{\mathrm{max}}+1}\cdot\exp\left(-j_{\mathrm{max}}\cdot\frac{x}{2\log_{2}(k)+1}\right)\geq p$, which leads to contradiction.
Now we assume $\tilde{j}$ satisfies (94). By the definition, we have that $y_{\tilde{j}}\geq q$.
By (93), we have that
$\mathsf{KL}(w^{\tilde{j}}\,\|\,w^{\tau_{\tilde{j}+1}})$ $\leq 2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta\varepsilon_{1}+(\tau_{{\tilde{j}}+1}-\tau_{\tilde{j}})\eta 2^{-(\tilde{j}-1)}$
$\quad+2\eta\sqrt{\frac{(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\log(1/\delta^{\prime})}{k}}+2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta^{2}+2\log(1/\delta^{\prime})\eta.$ (95)
Since $\log\left(\frac{y_{\tilde{j}+1}}{y_{\tilde{j}}}\right)\geq\frac{x}{2\log_{2}(k)+1}$ and $y_{\tilde{j}}\geq q$, by Lemma 23 and Lemma 24, we have that
@@
$\mathsf{KL}(w^{\tau_{\tilde{j}}}\,\|\,w^{\tau_{\tilde{j}+1}})\geq\mathsf{KL}\left(\mathrm{Ber}\left(y_{\tilde{j}}\right)\,\|\,\mathrm{Ber}\left(y_{\tilde{j}+1}\right)\right)\geq\frac{qx^{2}}{4(2\log_{2}(k)+1)^{2}},$
where $\mathrm{Ber}(x)$ denote the Bernoulli distribution with mean $x\in[0,1]$. As a result, we obtain that
$\frac{qx^{2}}{4(2\log_{2}(k)+1)^{2}}$ $\leq 2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta\varepsilon_{1}+(\tau_{{\tilde{j}}+1}-\tau_{\tilde{j}})\eta 2^{-({\tilde{j}}-1)}$
$\quad+2\eta\sqrt{\frac{(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\log(1/\delta^{\prime})}{k}}+2(\tau_{\tilde{j}+1}-\tau_{\tilde{j}})\eta^{2}+2\log(1/\delta^{\prime})\eta.$ (96)
By noting that $\frac{qx^{2}}{50(2\log_{2}(k)+1)^{2}}\geq\max\{2\log(1/\delta^{\prime})\eta,1/k\}$ and the fact that $\frac{2^{j_{\mathrm{max}}}}{\eta}\geq\frac{1}{\eta^{2}}$, we have that
$\tau_{\tilde{j}+1}-\tau_{\tilde{j}}$ $\geq\min\left\{\frac{qx^{2}}{100(2\log_{2}(k)+1)^{2}}\min\left\{\frac{1}{\eta\varepsilon_{1}},\frac{2^{\tilde{j}-1}}{\eta},\frac{1}{\eta^{2}}\right\},\frac{kx^{2}2^{-2l_{1}}}{10000\eta^{2}\log(1/\delta^{\prime})(\log_{2}(k)+1)^{4}}\right\}$
$\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}.$ (97)
By definition, we learn that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-j}]\geq\tau_{j+1}-t_{1}\geq\tau_{j+1}-\tau_{j}$ for any $11\leq j\leq 2\log_{2}(k)+$ 1. As a result we have that $\sum_{\tau=t_{1}}^{t_{2}-1}\mathbb{I}[-v^{\tau}\geq 2^{-\tilde{j}}]\geq\frac{qx^{2}\cdot 2^{\tilde{j}-1}}{100(2\log_{2}(k)+1)^{2}\eta}$.
The proof is completed by noting that $\tau_{j_{\mathrm{max}}+1}=t_{2}$.
Appendix D Missing proofs for learning Rademacher classes
D.1 Restatement and Proof of Lemma 15
Lemma 15. [Restatement] By running Algorithm 3, with probability $41-\delta/$ 4, it holds that
$L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ (98)
for any $1\leq t\leq T$.
Proof.
We will follow the notation in the proof of Lemma 4. Fix $n=\{n_{i}\}_{i=1}^{k}$ such that $n_{i}\geq 12\log(2k)$ for all $i\in[k]$ and $w\in\Delta(k)$. Let $\kappa=\min_{i}\frac{n_{i}}{w_{i}}$. Recall that $(x_{i,j},y_{i,j})$ is the $j$-th sample from $D_{i}$.
Define
$F(n,w):=\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)\right].$
Let $\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$ be a group of ghost samples for $\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}$. We then have that
$F(n,w)$ (99)
$\leq\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell(h,(x_{i,j},y_{i,j}))-\ell(h,(\tilde{x}_{i,j},\tilde{y}_{i,j}))\right)\right)\right]$
@@ D.1 Restatement and Proof of Lemma 15
$=\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{(\tilde{x}_{i,j},\tilde{y}_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\left(\ell(h,(x_{i,j},y_{i,j}))-\ell(h,(\tilde{x}_{i,j},\tilde{y}_{i,j}))\right)\right)\right]$
@@
$\leq 2\mathbb{E}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k},\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{1}{\kappa}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell(h,(x_{i,j},y_{i,j}))\right)\right]$ (100)
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}.$ (101)
Here (100) is by Lemma 30 as below.
Lemma 30.
Let $\mathcal{L}$ be a subset of $\mathbb{R}^{n}$. Let $w^{1},w^{2}\in\mathbb{R}^{n}$ be such that $|w^{1}_{i}|\leq|w_{i}^{2}|$ for all $i\in[n]$. Then it holds
$\mathbb{E}_{\{\sigma_{i}\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{1}_{i}f_{i}\right]\leq\mathbb{E}_{\{\sigma\}_{i=1}^{n}\overset{\mathrm{iid}}{\sim}\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w^{2}_{i}f_{i}\right].$ (102)
Proof.
It suffices to prove (102) under the case where $w_{i}^{1}=w^{2}_{i}$ for $11\leq i\leq n-$ 1, and $|w^{1}_{n}|\leq|w^{2}_{n}|$. Fix $\sigma_{i}$ for $11\leq i\leq n-$ 1.
$\mathbb{E}_{\sigma_{n}\sim\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}^{1}f_{i}\right]$
$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}f_{i}+w_{n}^{1}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}f_{i}-w_{n}^{1}f_{n}\right)$
$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{1}(f_{i}^{1}+f_{i}^{2})+w_{n}^{1}(f_{n}^{1}-f_{n}^{2})\right)$
$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+w_{n}^{1}(f_{n}^{1}-f_{n}^{2})\right)$
$\leq\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+|w_{n}^{2}(f_{n}^{1}-f_{n}^{2})|\right)$
$=\frac{1}{2}\max_{f^{1},f^{2}\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}(f_{i}^{1}+f_{i}^{2})+w_{n}^{2}(f_{n}^{1}-f_{n}^{2})\right)$
$=\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}f_{i}+w_{n}^{2}f_{n}\right)+\frac{1}{2}\max_{f\in\mathcal{L}}\left(\sum_{i=1}^{n-1}\sigma_{i}w_{i}^{2}f_{i}-w_{n}^{2}f_{n}\right)$
$=\mathbb{E}_{\sigma_{n}\sim\{\pm 1\}}\left[\max_{f\in\mathcal{L}}\sum_{i=1}^{n}\sigma_{i}w_{i}^{2}f_{i}\right].$
The proof is completed by taking expectation over $\{\sigma_{i}\}_{i=1}^{n-1}$. ∎
By Lemma 22 with $c=\frac{1}{\kappa}$, we have that
@@
$\mathrm{Pr}_{\{\{(x_{i,j},y_{i,j})\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\left|\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)-F(n,w)\right|\geq\varepsilon\right]$
$\leq$ $2\exp\left(-\frac{2\kappa^{2}\varepsilon^{2}}{\sum_{i=1}^{k}n_{i}}\right).$ (103)
By (101) and (103), for any $\delta^{\prime}\in(0,1]$, with probability $1-\delta^{\prime}$, it holds that
$\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right)$
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (104)
By replacing $(\ell,L)$ with $(-\ell,-L)$, using similar arguments we obtain that, with probability $1-2\delta^{\prime}$,
$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right|$
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (105)
Now we fix $\kappa\geq 0$. Define that
~=$\tilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\}.$
@@
By (105), for any $\delta^{\prime}>0$, with probability $1-\delta^{\prime}$, for any $\{n,w\}\in\tilde{\mathcal{L}}$, we have that
$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathbf{e}_{i})\right|$
@@
$\leq 2\frac{\sum_{i=1}^{k}n_{i}}{T_{1}}\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{T_{1}}\sqrt{\frac{d\log(2\sum_{i=1}^{k}n_{i})+\log(|\tilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}$
@@
$\leq 4\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{d\log(4T_{1})+\log(2|\tilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{T_{1}}}$
@@
$\leq 4\widetilde{\mathrm{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
@@
$\leq 600C_{T_{1}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}.$
Here the last inequality is by Assumption 10, Lemma 12, Lemma 13 and the fact that $\sum_{i=1}^{k}n_{i}\geq\frac{T_{1}}{2}$.
Note that in Algorithm 3,
$\hat{L}^{t}(h,w^{t})=\sum_{i=1}^{k}\frac{w_{i}^{t}}{\check{n}_{i}^{t}}\cdot\sum_{j=1}^{\check{n}_{i}^{t}}\ell(h,(x_{i,j},y_{i,j})).$
By the definition that $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ for $i\in[k]$, we have that $T_{1}w_{i}^{t}\leq\check{n}_{i}^{t}-1$ and $12\log(2k)\leq\check{n}_{i}^{t}\leq T_{1}$ for all $i\in[k]$. On the other hand, $\sum_{1=1}^{k}\check{n}_{i}^{t}\leq T_{1}+k+12k\log(2k)\leq 2T_{1}-2$. Therefore, there exists some $\tilde{w}^{t}\in\Delta(k)$ such that $\{\{\check{n}_{i}^{t}\}_{i=1}^{k},\tilde{w}^{t}\}\in\tilde{\mathcal{L}}$ and $\|w^{t}-\tilde{w}^{t}\|_{1}\leq\frac{\varepsilon_{1}}{8k}$ for each $1\leq t\leq T$. Choose $\delta^{\prime}=\delta/4$. We then obtain that: with probability $41-\delta/$ 4, it holds that
$\max_{h\in\mathcal{H}}|\hat{L}^{t}(h,w^{t})-L(h,w^{t})|$
$\leq 600C_{T_{1}}+4\sqrt{\frac{d\log(4T_{1})+2k\log(16kT_{1}/\epsilon_{1})+\log(8/\delta)}{T_{1}}}$
$\leq\frac{\varepsilon_{1}}{2}.$ (106)
for any $1\leq t\leq T$. Here the last inequality is by definition of $T_{1}$.
Finally, by noting that $h^{t}=\arg\min_{h\in\mathcal{H}}\hat{L}^{t}(h,w^{t})$, we have that
$L(h^{t},w^{t})\leq\hat{L}^{t}(h^{t},w^{t})+\frac{\varepsilon_{1}}{2}=\min_{h\in\mathcal{H}}\hat{L}^{t}(h,w^{t})+\frac{\varepsilon_{1}}{2}\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}.$
The proof is completed ∎
D.2 Restatement and Proof of Lemma 12
Lemma 12.[Restatement] For any two groups of positive integers $\{n_{i}\}_{i=1}^{k}$ and $\{m_{i}\}_{i=1}^{k}$, it holds that
Proof.

### v3

@@ D.4 Statement and proof of Lemma 20
Appendix E Proofs of auxiliary lemmas for Rademacher classes
E.1 Proof of Lemma 6
In this subsection, we shall follow the notation adopted in the proof of Lemma 1 (e.g., the dataset $\widetilde{\mathcal{S}}$, the data subset $\widetilde{\mathcal{S}}(n)$ and its independent copy $\widetilde{\mathcal{S}}^{+}(n)$, the Rademacher random variables $\{\sigma_{i}^{j}\}$).
Consider any given $n=\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for all $i\in[k]$, and any given $w\in\Delta(k)$. Let $\kappa=\min_{i}\frac{n_{i}}{w_{i}}$. Recall that $(x_{i,j},y_{i,j})$ is the $j$-th sample from $\mathcal{D}_{i}$, and $\{(x_{i,j}^{+},y_{i,j}^{+}\}$ are independent copies. Define
@@ E.1 Proof of Lemma 6
$F(n,w)$ $\coloneqq\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right)\right],$
which can be upper bounded by
$F(n,w)$ $=\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n)}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}^{+}(n)}\left[\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right]\right)\right)\right]$
$\leq\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n)}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\left(\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right)\right)\right]$
@@
$=\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n),\widetilde{\mathcal{S}}^{+}(n),\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\left\{\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\ell\big{(}h,(x_{i,j}^{+},y_{i,j}^{+})\big{)}\right\}\right)\right]$
@@
$\leq 2\mathop{\mathbb{E}}_{\widetilde{\mathcal{S}}(n),\{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}}\}_{i=1}^{k}}\left[\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{1}{\kappa}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}\right)\right]$
$=2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathrm{\mathsf{Rad}}}_{\{n_{i}\}_{i=1}^{k}}.$ (137)
Here, the first inequality arises from Jensen’s inequality, whereas the penultimate line applies Lemma 12.
Invoking the Mcdiarmid inequality (see Lemma 9) with the choice $c=1/\kappa$, we obtain
@@
$\mathbb{P}\left\{\left|\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\sum_{j=1}^{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right)-F(n,w)\right|\geq\varepsilon\right\}\leq 2\exp\left(-\frac{2\kappa^{2}\varepsilon^{2}}{\sum_{i=1}^{k}n_{i}}\right).$ (138)
@@
This taken together with (137) reveals that: for any $\delta^{\prime}\in(0,1]$, with probability at least $1-\delta^{\prime}$ we have
@@
$\max_{h\in\mathcal{H}}\left(\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell(h,(x_{i,j},y_{i,j}))-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right)\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (139)
Evidently, this inequality continues to hold if we replace $(\ell,L)$ with $(-\ell,-L)$. As a consequence, with probability at least $1-2\delta^{\prime}$ one has
@@
$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\frac{w_{i}}{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right|\leq 2\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{\kappa}\sqrt{\frac{\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}.$ (140)
Now, fix $\kappa\geq 0$, and define
~=$\widetilde{\mathcal{L}}=\left\{n=\{n_{i}\}_{i=1}^{k},w=\{w_{i}\}_{i=1}^{k}\in\Delta_{\varepsilon_{1}/(8k)}(k)\mid T_{1}w_{i}\leq 2n_{i},12\log(2k)\leq n_{i}\leq T_{1},\forall i\in[k],\sum_{i=1}^{k}n_{i}\leq 2T_{1}\right\},$
where $\Delta_{\varepsilon_{1}/(8k)}(k)$ (i.e., an $\varepsilon_{1}/(8k)$-net of $\Delta(k)$) has been defined in Appendix B.1. Inequality (140) combined with the union bound tells us that: for any $\delta^{\prime}>0$, with probability at least $1-\delta^{\prime}$
$\max_{h\in\mathcal{H}}\left|\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\frac{w_{i}}{n_{i}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}-\sum_{i=1}^{k}w_{i}L(h,\mathrm{e}^{\mathsf{basis}}_{i})\right|$
@@
$\qquad\leq\frac{\sum_{i=1}^{k}n_{i}}{T_{1}/2}\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+\frac{\sum_{i=1}^{k}n_{i}}{T_{1}/2}\sqrt{\frac{\log(|\widetilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{2\sum_{i=1}^{k}n_{i}}}$
@@
$\qquad\leq 4\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{\log(2|\widetilde{\mathcal{L}}|)+\log(2/\delta^{\prime})}{T_{1}}}$
$\qquad\leq 4\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}+4\sqrt{\frac{2k\log(16kT_{1}/\varepsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
$\qquad\leq 600C_{T_{1}}+4\sqrt{\frac{2k\log(16kT_{1}/\varepsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}$
holds simultaneously for all $\{n,w\}\in\widetilde{\mathcal{L}}$; see the use of the union bound in Appendix B.1 too. Here, we have made use of Assumption 1, Lemma 4, Lemma 5 and the fact that $\sum_{i=1}^{k}n_{i}\geq T_{1}/2$.
Note that in Algorithm 3, we take
$\widehat{L}^{t}(h,w^{t})=\sum_{i=1}^{k}\frac{w_{i}^{t}}{n_{i}^{t,\mathsf{rad}}}\cdot\sum_{j=1}^{n_{i}^{t,\mathsf{rad}}}\ell(h,(x_{i,j},y_{i,j})).$
Given our choice that $n_{i}^{t,\mathsf{rad}}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ for $i\in[k]$, we see that
$T_{1}w_{i}^{t}\leq n_{i}^{t,\mathsf{rad}}-1\qquad\text{and}\qquad 12\log(2k)\leq n_{i}^{t,\mathsf{rad}}\leq T_{1}$
for all $i\in[k]$. In addition, it is seen from our choice of $n_{i}^{t,\mathsf{rad}}$ and $T_{1}$ that
$\sum_{i=1}^{k}n_{i}^{t,\mathsf{rad}}\leq\sum_{i=1}^{k}\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil\leq T_{1}+k+12k\log(2k)\leq 2T_{1}-2.$
Therefore, there exists some $\widetilde{w}^{t}\in\Delta(k)$ satisfying
$\big{\{}\{n_{i}^{t,\mathsf{rad}}\}_{i=1}^{k},\widetilde{w}^{t}\big{\}}\in\widetilde{\mathcal{L}}\qquad\text{and}\qquad\|w^{t}-\widetilde{w}^{t}\|_{1}\leq\frac{\varepsilon_{1}}{8k}$
for each $1\leq t\leq T$. Taking $\delta^{\prime}=\delta/4$, we obtain that with probability at least $41-\delta/$ 4,
$\max_{h\in\mathcal{H}}\big{|}\widehat{L}^{t}(h,w^{t})-L(h,w^{t})\big{|}$ $\leq\max_{h\in\mathcal{H}}\big{|}\widehat{L}^{t}(h,\widetilde{w}^{t})-L(h,\widetilde{w}^{t})\big{|}+\max_{h\in\mathcal{H}}\big{|}\widehat{L}^{t}(h,\widetilde{w}^{t})-\widehat{L}^{t}(h,w^{t})\big{|}$
$\qquad\qquad+\max_{h\in\mathcal{H}}\big{|}L(h,\widetilde{w}^{t})-L(h,w^{t})\big{|}$
$\leq 600C_{T_{1}}+4\sqrt{\frac{2k\log(16kT_{1}/\varepsilon_{1})+\log(2/\delta^{\prime})}{T_{1}}}+\frac{\varepsilon_{1}}{8k}+\frac{\varepsilon_{1}}{8k}$
$\leq\frac{\varepsilon_{1}}{2}$
for any $1\leq t\leq T$, where the last inequality results from the definition of $T_{1}$.
Finally, the fact that $h^{t}=\arg\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})$ allows one to derive
$L(h^{t},w^{t})\leq\widehat{L}^{t}(h^{t},w^{t})+\frac{\varepsilon_{1}}{2}=\min_{h\in\mathcal{H}}\widehat{L}^{t}(h,w^{t})+\frac{\varepsilon_{1}}{2}\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1},$
which concludes the proof.
E.2 Proof of Lemma 4